# Gaussian-Forest: Hierarchical-Hybrid 3D Gaussian Splatting for Compressed Scene Modeling

### v1

@@ 3.2 Gaussian-Forest Modeling
Hybrid Tree Nodes. The architecture consists of three distinct node types: leaf node $\mathbf{T}_{\text{E}}$ encapsulating explicit attributes, internal node $\mathbf{T}_{\text{I}}$, and root node $\mathbf{T}_{\text{R}}$, both dedicated to implicit attribute modeling. Each node type is characterized by a unique set of parameters:
@@
$\begin{split}&\mathbf{T}_{\text{E}}^{(2)}:=\{\bm{\mu},\gamma_{s},\alpha,p^{(2)}\}\in\mathbb{R}^{6},\\ &\mathbf{T}_{\text{I}}^{(1)}:=\{\mathbf{f}_{\text{I}},p^{(1)}\}\in\mathbb{R}^{\mathcal{D}_{I}+1},\mathbf{T}^{(0)}_{\text{R}}:=\{\mathbf{f}_{\text{R}}\}\in\mathbb{R}^{\mathcal{D}_{R}},\end{split}$ (3)
@@
with the superscript indicating the node depth within the tree. Among these notations, $\gamma_{s}$ functions as the scaling coefficient for the implicit scaling $\mathbf{s}$, $p$ represents an integer index pointer, and $\mathbf{f}$ signifies a feature vector, with further elaboration following. The rationale beyond this hybrid modeling of Gaussian parameters lies in the fact that explicit attributes like position $\bm{\mu}$ and opacity $\alpha$ are agnostic to the viewing angle. Conversely, view-dependant attributes like covariance-related rotation $\mathbf{q}$, scaling $\mathbf{s}$, and colors $\mathbf{c}$ are prone to redundancy. This motivates us to model them implicitly and designate significantly fewer nodes at higher levels, i.e., $N_{\text{R}}\ll N_{\text{I}}\ll N$, where $N_{\text{R}}$ and $N_{\text{I}}$ represents the number of root and internal nodes, respectively. All these number are adaptively adjusted during the forest’s growth and pruning processes (Sec. 3.3).
@@
Leaf Node Traversal. Under the formulation of Eq. (3), the pointer $p^{(l)}$ stored at node $\mathbf{T}^{(l)}$ in the $l$-th layer establishes a link from $\mathbf{T}^{(l)}$ to its parent node at the ($l-1$)-th layer. By recursively tracing pointers upwards, each leaf node $\mathbf{T}_{\text{E}}^{(2)}$ traverses a unique path $[\mathbf{T}_{\text{E}}^{(2)},\mathbf{T}_{\text{I}}^{(1)},\mathbf{T}_{\text{R}}^{(0)}]$ leading to the root $\mathbf{T}_{\text{R}}^{(0)}$ of its corresponding tree, collectively defining the parameters of a distinct Gaussian. For the representation of implicit attributes, we concatenate the latent features along this path as $\mathbf{f}=[\mathbf{f}_{\text{I}},\mathbf{f}_{\text{R}}]\in\mathbb{R}^{\mathcal{D}_{I}+\mathcal{D}_{R}}$.
@@
View-dependent Implicit Attributes. To model the implicit attributes, we employ two MLP $\mathcal{F}_{\text{cov}}$ and $\mathcal{F}_{\text{rgb}}$ to decode the obtained latent features $\mathbf{f}$. This decoding results in: (1) the covariance-related scaling $\mathbf{s}$ and rotation vectors $\mathbf{q}$, and (2) the view-dependent color $\mathbf{c}$, as shown below:
@@
This hierarchical representation offers efficiency advantages by reducing over 80% of parameters while retaining adaptability and expressive capabilities. Detailed theoretical and empirical analyses regarding storage size are presented in Sec. 4.4.3. In conclusion, Gaussian parameterization in Eq. (2) can be rewritten as:
@@
This formulation minimizes the number of parameters for each Gaussian component. However, this minimization is effective only when the total number of nodes is controlled. In the subsequent section, we elaborate on strategies for optimizing the Gaussian-Forest structure.
Table 1: Quantitative comparisons on Mip-NeRF360, Tanks&Temples, Deep Blending and Synthetic Blender datasets. The best and second-best outcomes are shown in bold deep blue and light blue, respectively. All scores for compared methods are sourced from published papers or released pre-trained models, except for the hyphen (-) indicating no valid data and signifying re-evaluation on our machine.
@@
FPS MB FPS/MB SSIM PSNR Train FPS MB FPS/MB SSIM PSNR Train
Plenoxels Fridovich-Keil et al. (2022) 6.79 2150 0.003 0.626 23.08 26 m 13.0 2355 0.006 0.719 21.08 25 m
NGP-Base Müller et al. (2022) 11.7 13 0.900 0.671 25.30 6 m 17.1 13 1.315 0.723 21.72 5 m
NGP-Big Müller et al. (2022) 9.43 48 0.196 0.699 25.59 8 m 14.4 48 0.300 0.745 21.92 7 m
Mip-360 Barron et al. (2022) 0.06 8.6 0.007 0.792 27.69 48 h 0.14 8.6 0.016 0.759 22.22 48 h
3DGS Kerbl et al. (2023) 134 734 0.183 0.815 27.21 42 m 154 411 0.375 0.841 23.14 27 m
3DGS Kerbl et al. (2023) 105 827 0.127 0.816 27.45 28 m 143 454 0.315 0.848 23.73 16 m
GaussianForest-Large 105 85 1.235 0.803 27.45 28 m 164 45 3.644 0.839 23.67 16 m
GaussianForest-Small 121 50 2.426 0.797 27.33 26 m 175 38 4.605 0.836 23.56 15 m
@@
FPS MB FPS/MB SSIM PSNR Train FPS MB FPS/MB SSIM PSNR Train
Plenoxels Fridovich-Keil et al. (2022) 11.2 2765 0.004 0.795 23.06 28 m - 778 - 0.958 31.71 11 m
NGP-Base Müller et al. (2022) 3.26 13 0.251 0.797 23.62 7 m - 13 - 0.963 33.18 5 m
NGP-Big Müller et al. (2022) 2.79 48 0.058 0.817 24.96 8 m - - - - - -
Mip-360 Barron et al. (2022) 0.09 8.6 0.010 0.901 29.40 48 h - 8.6 - 0.961 33.09 48 h
3DGS Kerbl et al. (2023) 137 676 0.203 0.903 29.41 36 m - - - - 33.32 -
3DGS Kerbl et al. (2023) 106 701 0.151 0.904 29.54 25 m 344 72 4.778 0.969 33.80 7 m
GaussianForest-Large 96 98 0.980 0.908 30.18 29 m 417 11 37.91 0.969 33.60 7 m
GaussianForest-Small 107 64 1.672 0.905 30.11 25 m 445 8.5 52.71 0.967 33.52 6 m
3.3 Forest Growing and Pruning
@@
To control the number of nodes required for scene modeling, Gaussian-Forest is initialized as a small set of singly linked lists and undergoes adaptive growth and pruning to evolve into an efficient and robust forest. Specifically, branches exhibiting underfitting or high uncertainty are selectively expanded by adding more leaf and/or non-leaf nodes. In order to prevent excessive growth of the forest, we implement early stopping and pruning strategies. These approaches are pivotal in maintaining a concise yet faithful presentation, accelerating both the training and rendering processes.
3.3.1 Initialization
@@
With the given SfM point cloud containing $N_{\text{SfM}}$ points, we correspondingly establish $N_{\text{SfM}}$ leaf nodes with explicit attributes initialized according to 3DGS Kerbl et al. (2023). Following this, we initialize the root and internal layer, with each layer comprising $K$ nodes where $K\ll N_{\text{SfM}}$. Nodes in two consecutive layers are interconnected in a one-to-one manner. Subsequently, we employ the K-means algorithm to group the leaf nodes into $K$ clusters based on proximity. Each leaf node is then connected to an internal node corresponding to its cluster, thereby forming $N_{\text{SfM}}$ singly linked lists. All implicit feature vectors in $\mathbf{T}_{\text{I}}^{(1)}$ and $\mathbf{T}_{\text{R}}^{(0)}$ are randomly initialized. For scenes like Synthetic Blending Mildenhall et al. (2020) with no available SfM point clouds, leaf nodes are initialized using $N_{\text{SfM}}=100$k synthetic points generated by uniform sampling following 3DGS Kerbl et al. (2023).
3.3.2 Forest Growth
@@
To adapt to varying complexities in different scenes, a hierarchical forest growth strategy is leveraged based on cumulative gradients $\mathbf{CG}$ of leaf nodes during the end-to-end optimization. These gradients serve as indicators of the learning difficulty for each of the $N$ Gaussians. Specifically, we consider three distinct cases governed by a set of gradient thresholds denoted by $\{\mathcal{T}_{l}\}_{l\in[0,L)}$ arranged in non-increasing order, while nodes beyond these three cases remain unchanged.
@@
Case 0: $\mathcal{T}_{2}<\mathbf{CG}\leq\mathcal{T}_{1}$. For leaf nodes satisfying this case, growth is limited to their own cloning, creating a new link to each of their original parent nodes. This aligns with the split strategy in 3DGS Kerbl et al. (2023).
@@
All these cases are illustrated in Figure 2. The motivation behind this hierarchical design is that a minimal cumulative gradient of a Gaussian implies the sufficient representational capacity for its local area. Hence, simpler background regions can be effectively modeled with fewer Gaussians. Conversely, a high cumulative gradient indicates the need for more detailed features, especially in complex regions like object boundaries or areas varying from different angles. To address this, both the leaves and their parent nodes are replicated, and the original leaves and their clones are then linked to the newly formed non-leaf nodes, thereby enhancing the model’s ability to depict these intricate areas with finer detail through increased feature dimensions.
3.3.3 Early Stopping
@@ 3.3.4 Forest Pruning
4 Experiments
4.1 Implementation Details
Framework and Hardware. Our Gaussian-Forest is implemented based on 3DGS Kerbl et al. (2023) and PyTorch. All experiments were conducted on a GeForce RTX 3090 GPU, which shares the same CUDA compute capability (8.6) as the RTX A6000 GPU used in 3DGS Kerbl et al. (2023). For fair comparison, we re-executed their code on our machine.
Forest Structure. We instantiate the Gaussian-Forest as a composition of trees with $L=3$ layers: one root layer, one internal layer, and one leaf layer. Each implicit layer is initialized with $K=10$k nodes, and the feature dimensions for each layer are specified as $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=\{24,16\}$ and $\{32,24\}$ for the Large and Small settings, respectively.
Forest Growth. Forest growth occurs every 100 iterations, with growth thresholds set at $\{\mathcal{T}_{l}\}=\{1\times 10^{-3},2.5\times 10^{-4},2\times 10^{-4}\}$. The number of iterations to stop the growth of each layer is defined as $\{t_{l}\}$ = {5k, 10k, 15k}, and the training process concludes after the 30k iterations.
@@ 4.1 Implementation Details
Figure 3: Visualization of quantitative comparisons on Mip-NeRF360 and Tanks & Temples datasets. The horizontal and vertical axes represent rendering speed and quality, respectively. Each point’s size in the figure indicates the corresponding model size in MB.
@@
Forest Pruning. Gaussians with $\alpha<\mathcal{T}_{\alpha}$ or $\gamma_{s}<\mathcal{T}_{s}$ are identified and pruned every 100 iterations, where $\{\mathcal{T}_{\alpha},\mathcal{T}_{s}\}=\{1\times 10^{-2},5\times 10^{-4}\}$. In contrast to the early-stop strategy for forest growth, pruning continues until the end of training, with a larger interval defined as 1,000 iterations.

### v2

@@ III-B1 Hybrid Tree Nodes
The architecture consists of three distinct node types: leaf node $\mathbf{T}_{\text{L}}$ encapsulating explicit attributes; internal node $\mathbf{T}_{\text{I}}$ and root node $\mathbf{T}_{\text{R}}$, both dedicated to implicit attribute modeling. Each node type is characterized by a unique set of parameters:
@@
$\begin{split}&\mathbf{T}_{\text{L}}^{(2)}:=\{\bm{\mu},\gamma_{s},\alpha,p^{(2)}\}\in\mathbb{R}^{6},\\ &\mathbf{T}_{\text{I}}^{(1)}:=\{\mathbf{f}_{\text{I}},p^{(1)}\}\in\mathbb{R}^{\mathcal{D}_{I}+1},\\ &\mathbf{T}^{(0)}_{\text{R}}:=\{\mathbf{f}_{\text{R}}\}\in\mathbb{R}^{\mathcal{D}_{R}},\end{split}$ (3)
with the superscript indicating the depth within the tree. Among these notations, $\gamma_{s}$ functions as a scaling coefficient for the implicit scaling $\mathbf{s}$, $p$ represents an integer index pointer, and $\mathbf{f}$ signifies a feature vector, all of which will be further elaborated upon in the following. Note that such a structure can easily be extended by adding more internal layers.
@@
The rationale behind this hybrid modeling of Gaussian parameters lies in the observation that attributes like position $\bm{\mu}$ and opacity $\alpha$ exhibit sensitivity to deviations, are invariant to viewing angle, and can vary significantly across neighbors. Consequently, we classify these as explicit attributes and model them directly in the leaf nodes. In contrast, attributes like covariance-related rotation $\mathbf{q}$, scaling $\mathbf{s}$, and view-dependent colors $\mathbf{c}$ encompass more parameters but typically exhibit local smoothness, leading to redundancy when modeled explicitly. This insight motivates us to model these attributes implicitly, assigning significantly fewer nodes at higher hierarchical levels, i.e., $N_{\text{R}}\ll N_{\text{I}}\ll N$, where $N_{\text{R}}$ and $N_{\text{I}}$ denote the number of root and internal nodes, respectively. All these number are adaptively adjusted during the forest’s growth and pruning processes (Sec. III-C).
III-B2 Node Traversal
@@
Under the formulation of Eq. (3), the pointer $p^{(l)}$ stored at node $\mathbf{T}^{(l)}$ in the $l$-th layer establishes a link from $\mathbf{T}^{(l)}$ to its parent node at the ($l-1$)-th layer. By recursively tracing pointers upwards, each leaf node $\mathbf{T}_{\text{L}}^{(2)}$ traverses a unique path $[\mathbf{T}_{\text{L}}^{(2)},\mathbf{T}_{\text{I}}^{(1)},\mathbf{T}_{\text{R}}^{(0)}]$ leading to the root $\mathbf{T}_{\text{R}}^{(0)}$ of its corresponding tree, collectively defining the parameters of a distinct 3D Gaussian. For the representation of implicit attributes, we concatenate the latent features along this path as $\mathbf{f}=[\mathbf{f}_{\text{I}},\mathbf{f}_{\text{R}}]\in\mathbb{R}^{\mathcal{D}_{I}+\mathcal{D}_{R}}$.
III-B3 Implicit Attributes
@@
To model the implicit attributes, we employ two MLP $\mathcal{F}_{\text{cov}}$ and $\mathcal{F}_{\text{rgb}}$ to decode the obtained latent features $\mathbf{f}$. This decoding results in: (1) the covariance-related scaling $\mathbf{s}$ and rotation vectors $\mathbf{q}$, and (2) the view-dependent color $\mathbf{c}$, as shown below:
@@
This hierarchical representation offers efficiency advantages by reducing over 80% of parameters while retaining adaptability and expressive capabilities. Detailed theoretical and empirical analyses regarding storage size are presented in Sec. IV-D3. In conclusion, Gaussian parameterization in Eq. (2) can be rewritten as:
@@
This formulation minimizes the number of parameters for each Gaussian component. However, this minimization is effective only when the total number of nodes is controlled. In the subsequent section, we elaborate on strategies for optimizing the GaussianForest structure.
III-B4 Motivation Illustration
As previously mentioned, our motivation arises from the high similarity among local Gaussians, which results in substantial parameter redundancy when modeled individually, as shown in 3DGS [2]. Fig. 3 exemplifies this insight and intuitively demonstrates the efficiency and potential of GaussianForest. In this figure, we visualize two Gaussian trees by reducing the opacity of the others for clear observation. Specifically, we maintain the opacity of all Gaussians associated with two selected roots while reducing the opacity of the remaining ones to one-tenth.
Figure 3: Visualization of two Gaussian trees. A drum head is approximately modeled by a single Gaussian tree, requiring far fewer parameters compared to the thousands of explicit Gaussians in 3DGS. GaussianForest also shows inherent clustering ability, naturally segmenting similar regions without any supervisory information. During optimization, adjacent regions with similar geometric and color features tend to aggregate under the same parent node.
As illustrated in Fig. 3, one Gaussian tree effectively models a drum head, while the other represents a portion of a cymbal. It is noteworthy that without this hierarchical-hybrid design, modeling a drum head would necessitate thousands of Gaussians, each with 59 free parameters. With GaussianForest, the same task is accomplished using only one root and dozens of internal nodes, along with lightweight leaves with only 6 parameters, thereby significantly reducing the number of required parameters.
Another interesting observation from Fig. 3 is that Gaussians belonging to the same parent node exhibit similar positional, geometric, and color characteristics. In other words, during the optimization of the GaussianForest, adjacent regions with similar geometric and color features tend to aggregate under the same parent node. This inherent clustering ability naturally segments similar regions without any supervisory information. As illustrated, GaussianForest approximately delineates a whole drum head, from which we observe the potential of leveraging GaussianForest for unsupervised 3D segmentation and scene understanding.
TABLE I: Quantitative comparisons on Mip-NeRF360, Tanks&Temples, Deep Blending and Synthetic Blender datasets. The best and second-best outcomes are shown in bold deep blue and light blue, respectively. All scores for compared methods are sourced from published papers or released pre-trained models, except for the hyphen (-) indicating no valid data and signifying re-evaluation on our machine.
@@ III-B4 Motivation Illustration
Train FPS Size FPS/MB SSIM PSNR LPIPS Train FPS MB FPS/MB SSIM PSNR LPIPS
Plenoxels [20] 26 m 6.79 2150 0.003 0.626 23.08 0.463 25 m 13.0 2355 0.006 0.719 21.08 0.379
NGP-Base [24] 6 m 11.7 13 0.900 0.671 25.30 0.371 5 m 17.1 13 1.315 0.723 21.72 0.330
NGP-Big [24] 8 m 9.43 48 0.196 0.699 25.59 0.331 7 m 14.4 48 0.300 0.745 21.92 0.305
Mip-360 [28] 48 h 0.06 8.6 0.007 0.792 27.69 0.237 48 h 0.14 8.6 0.016 0.759 22.22 0.257
3DGS [2] 42 m 134 734 0.183 0.815 27.21 0.214 27 m 154 411 0.375 0.841 23.14 0.183
3DGS [2] 28 m 105 827 0.127 0.816 27.45 0.201 16 m 143 454 0.315 0.848 23.73 0.179
GF-Large 28 m 105 85 1.235 0.803 27.45 0.212 16 m 164 45 3.644 0.839 23.67 0.188
GF-Small 26 m 121 50 2.426 0.797 27.33 0.219 15 m 175 38 4.605 0.836 23.56 0.194
@@
Train FPS MB FPS/MB SSIM PSNR LPIPS Train FPS MB FPS/MB SSIM PSNR LPIPS
Plenoxels [20] 28 m 11.2 2765 0.004 0.795 23.06 0.510 11 m - 778 - 0.958 31.71 -
NGP-Base [24] 7 m 3.26 13 0.251 0.797 23.62 0.423 5 m - 13 - 0.963 33.18 -
NGP-Big [24] 8 m 2.79 48 0.058 0.817 24.96 0.390 - - - - - - -
Mip-360 [28] 48 h 0.09 8.6 0.010 0.901 29.40 0.245 48 h - 8.6 - 0.961 33.09 -
3DGS [2] 36 m 137 676 0.203 0.903 29.41 0.243 - - - - - 33.32 -
3DGS [2] 25 m 106 701 0.151 0.904 29.54 0.221 7 m 344 72 4.778 0.969 33.80 0.002
GF-Large 29 m 96 98 0.980 0.908 30.18 0.215 7 m 417 11 37.91 0.969 33.60 0.002
GF-Small 25 m 107 64 1.672 0.905 30.11 0.223 6 m 445 8.5 52.71 0.967 33.52 0.002
III-C Forest Growing and Pruning
@@
To control the number of nodes required for scene modeling, GaussianForest is initialized as a small set of singly linked lists and undergoes adaptive growth and pruning to evolve into an efficient and robust forest. Specifically, branches exhibiting underfitting or high uncertainty are selectively expanded by adding more leaf and/or non-leaf nodes. In order to prevent excessive growth of the forest, we implement early stopping and pruning strategies. These approaches are pivotal in maintaining a concise yet faithful presentation, accelerating both the training and rendering processes.
III-C1 Initialization
@@
With the given SfM point cloud containing $N_{\text{SfM}}$ points, we correspondingly establish $N_{\text{SfM}}$ leaf nodes with explicit attributes initialized according to 3DGS [2]. Following this, we initialize the root and internal layer, with each layer comprising $K$ nodes where $K\ll N_{\text{SfM}}$. Nodes in two consecutive layers are interconnected in a one-to-one manner. Subsequently, we employ the K-means algorithm to group the leaf nodes into $K$ clusters based on proximity. Each leaf node is then connected to an internal node corresponding to its cluster, thereby forming $N_{\text{SfM}}$ singly linked lists. All implicit feature vectors in $\mathbf{T}_{\text{I}}^{(1)}$ and $\mathbf{T}_{\text{R}}^{(0)}$ are randomly initialized. For scenes like Synthetic Blending [1] with no available SfM point clouds, leaf nodes are initialized using $N_{\text{SfM}}=100$k synthetic points generated by uniform sampling following 3DGS [2].
III-C2 Forest Growth
@@
To adapt to varying complexities in different scenes, a hierarchical forest growth strategy is leveraged based on cumulative gradients $\mathbf{CG}$ of leaf nodes during the end-to-end optimization. These gradients serve as indicators of the learning difficulty for each Gaussians. Specifically, we consider three distinct cases governed by a set of gradient thresholds denoted by $\{\mathcal{T}_{l}\}_{l\in[0,L)}$ arranged in non-increasing order, while nodes beyond these cases remain unchanged.
@@
Case 0: $\mathcal{T}_{2}<\mathbf{CG}\leq\mathcal{T}_{1}$. For leaf nodes satisfying this case, growth is limited to their own cloning, creating a new link to each of their original parent nodes. This aligns with the split strategy in 3DGS [2].
@@
All these cases are illustrated in Figure 2. The motivation behind this hierarchical design is that a minimal $\mathbf{CG}$ of a Gaussian implies the sufficient representational capacity for its local area. Hence, simpler background regions can be effectively modeled with fewer Gaussians. Conversely, a high $\mathbf{CG}$ indicates the need for more detailed features, especially in complex regions like object boundaries or areas varying from different angles. To address this, both the leaves and their parent nodes are replicated, and the original leaves and their clones are then linked to the newly formed non-leaf nodes, thereby enhancing the model’s ability to depict these intricate areas with finer detail through increased feature dimensions.
III-C3 Early Stopping
@@ III-C4 Forest Pruning
Figure 4: Visualization of quantitative comparisons on Mip-NeRF360 and Tanks&Temples datasets. The horizontal and vertical axes represent rendering speed and quality, respectively. Each point’s size in the figure indicates the corresponding model size in MB. This comparison serves to highlight the superiority of our approach.
IV Experiments
IV-A Implementation Details
IV-A1 Framework and Hardware
Our GaussianForest is implemented based on 3DGS [2] and PyTorch. All experiments were conducted on a GeForce RTX 3090 GPU, which shares the same CUDA compute capability (8.6) as the RTX A6000 GPU used in 3DGS [2]. For fair comparison, we re-executed their code on our machine.
IV-A2 Forest Structure
We instantiate the GaussianForest as a composition of trees with $L=3$ layers: one root layer, one internal layer, and one leaf layer. Each implicit layer is initialized with $K=10$k nodes, and the feature dimensions for each layer are specified as $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=\{24,16\}$ and $\{32,24\}$ for the Small and Large settings, respectively.
IV-A3 Forest Growth
Forest growth occurs every 100 iterations, with growth thresholds set at $\{\mathcal{T}_{l}\}=\{1\times 10^{-3},2.5\times 10^{-4},2\times 10^{-4}\}$. The number of iterations to stop the growth of each layer is defined as $\{t_{l}\}$ = {5k, 10k, 15k}, and the training process concludes after the 30k iterations.
IV-A4 Forest Pruning
@@
Gaussians with $\alpha<\mathcal{T}_{\alpha}$ or $\gamma_{s}<\mathcal{T}_{s}$ are identified and pruned every 100 iterations, where $\{\mathcal{T}_{\alpha},\mathcal{T}_{s}\}=\{1\times 10^{-2},5\times 10^{-4}\}$. In contrast to the early-stop strategy for forest growth, pruning continues until the end of training, with a larger interval defined as 1,000 iterations.
IV-A5 Features and Decoders